perform the indicated operations
9/16 - 1/12

Answer by Alan3354(69443)

9/16 - 1/12
--------------
Find a common denominator. The product always works.
16*12 = 192
------------
9/16 - 1/12
108/192 - 16/192 = 92/192
= 23/48
------------
You can waste time finding the LCD, least common denominator, but if it's not asked for, it only wastes time. That can make a difference on a test.